A point moves so that the angle from the line joining it and the origin to the line (3, -2) and (5, 7) is 45º. Find the equation of the locus.



1
Expert's answer
2021-10-18T15:16:24-0400

Solution:

Let the coordinates of moving point be P(h,k).

The equation of line joining P(h,k) and origin (0,0) is:

"y-k=\dfrac{k-0}{h-0}(x-h)
\\\Rightarrow y-k=\dfrac{k}{h}(x-h)
\\\Rightarrow y-k=\dfrac{kx}h-k
\\\Rightarrow y=\dfrac{kx}h \ ...(i)"

Here, slope is "m_1=\dfrac kh"

Now, the equation of line joining (3,-2), (5,7) is:

"y-(-2)=\dfrac{7+2}{5-3}(x-3)
\\\Rightarrow y+2=\dfrac{9}{2}(x-3)
\\\Rightarrow y+2=\dfrac{9x}{2}-\dfrac{27}{2}
\\\Rightarrow y=\dfrac{9x}{2}-\dfrac{31}{2}\ ...(ii)"

Here, slope is "m_2=\dfrac92"

Angle between the lines is "\theta":

"\tan \theta=|\dfrac{m_1-m_2}{1+m_1m_2}|
\\\Rightarrow \tan 45\degree=|\dfrac{\dfrac kh-\dfrac 92}{1+\dfrac kh\times \dfrac 92}|"

"\\\Rightarrow 1=|\dfrac{\dfrac {2k-9h}{2h}}{\dfrac {2h+9k}{2h}}|
\\\Rightarrow \pm1=\dfrac {2k-9h}{2h+9k}"

"\\\Rightarrow 2h+9k=2k-9h;\ \ 2h+9k=9h-2k
\\\Rightarrow 11h+7k=0; \ \ 7h-11k=0
\\\Rightarrow 11x+7y=0; \ \ 7x-11y=0"